A recent college graduate got a good job and began a savings account. He authorized the bank to automatically transfer $500 each month from his checking account to the savings account. The bank made the first withdrawal on July 1, 2015 and is instructed to make the last withdrawal on January 1, 2040. The bank pays a nominal interest rate of 6% and compounds twice a month. What is the future worth of the account on January 1, 2040?
Answer #1

Monthly withdrawals = $ 500, Tenure of Withdrawals = July1, 2015 to January 1, 2040 or (25.5 years)

Tenure in Months = 25.5 x 12 = 306 months

Nominal Interest Rate = 6 % with compounding twice a month

Effective Monthly Interest Rate = [1+(0.06/24)]^(2) - 1 = 0.005006 or 0.5006 %

Therefore, Future Worth = 500 x (1.005006)^(306) + 500 x (1.005006)^(305) +........+ 500 x (1.005006)^(1) = 500 x [{(1.005006)^(306) -1} / {1.005006 - 1}] = $ 3604691.